Show that the sequence {𝑓𝑛} where 𝑓𝑛(π‘₯) = 𝑛π‘₯(1 βˆ’ π‘₯)

𝑛 does not converge uniformly on [0, 1].


1
Expert's answer
2021-03-08T18:44:41-0500

Show that the sequence {𝑓𝑛} where 𝑓𝑛(π‘₯) = 𝑛π‘₯(1 βˆ’ π‘₯)𝑛 

does not converge uniformly on [0, 1].

Solution. Find the limit function "f(x)=\lim\limits_{n\to\infty}f_n(x)=0," where "\lim\limits_{n\to\infty}nx(1-x)^n=\lim\limits_{n\to\infty}\frac{nx}{(1-x)^{-n}}=[\frac{\infty}{\infty}]=-\lim\limits_{n\to\infty}\frac{x(1-x)^{n}}{\ln(1-x)}=0"

is found by Lopital-Bernoulli rule. According to the criterion of uniform

convergence, functional sequence "f_n(x)=nx(1-x)^n" does not

converge uniformly on [0,1]. Really, "\lim\limits_{n\to\infty}\sup\limits_{0\le x\le 1}|f(x)-f_n(x)|="

"\lim\limits_{n\to\infty}\sup\limits_{0\le x\le 1}(nx(1-x)^n)=" "\lim\limits_{n\to\infty}(1-\frac{1}{n+1})^{n+1}=\frac{1}{e}\ne 0."

Answer: "f_n(x)=nx(1-x)^n" does not converge uniformly on [0,1].